Find two non negative integers that have the same answer when multiplied together as when one is added by the other
Answer by Edwin McCravy(20086)

Two solutions, 

1. the pair of non-negative integers 0 and 0

0×0 = 0
0+0 = 0

2. the pair of non-negative integers 2 and 2

2×2 = 4
2+2 = 4

Edwin